OpenStudy (anonymous):

4+5=18 7+4=30 9+3=24 6+3=???? there is a logic to this that should reveal the answer pls help

OpenStudy (shamil98):

Those are all wrong. 4 + 5 = 9 7 + 4 = 11 9 + 3 = 12 6 + 3 = 9 o.o

OpenStudy (anonymous):

what do you mean?

OpenStudy (usukidoll):

troll attempt detected

OpenStudy (anonymous):

its a riddle someone gave me

OpenStudy (usukidoll):

@UnkleRhaukus can you help me with laplace? ^^